OpenStudy (anonymous):

Find the value of y so that the line passing through (6, y) and (7, 5) has a slope of -3

jimthompson5910 (jim_thompson5910):

m = (y2 - y1)/(x2 - x1) -3 = (5 - y)/(7 - 6) -3 = (5 - y)/1 -3 = 5 - y Keep going to solve for y

OpenStudy (anonymous):

find the value of y so that the line passes through the point ( -2,1) and (4,y) has a sloope of 2/3